N and M Corp. is considering leasing a new machine for ​$25000 comma per year. The lease arrangement calls for a 5​-year lease with an option to purchase the machine at the end of the lease for ​$5000. The firm is in the 26​% tax bracket. What is the present value of the lease​ outflows, including the purchase​ option, if lease payments are made at the end of each year and if the​ after-tax cost of debt is 8​%?
Answer #1

Post-tax cash flow = 25,000 x (1 - 26%) = 18,500

PV = CF1 / (1 + r) + CF2 / (1 + r)^2 + CF3 / (1 + r)^3 + CF4 / (1 + r)^4 + CF5 / (1 + r)^5

= 18,500 / 1.08 + 18,500 / 1.08^2 + ... + (18,500 + 5,000) / 1.08^5

= $77,268